OpenStudy (anonymous):

find the integral of (x^2)(1-x)^1/2, i get that you times it with 1/3 and you set 1-x to zero and get answer then plug that in, but i am lost with the formula help please

OpenStudy (zarkon):

u=1-x

OpenStudy (anonymous):

yes and du is \[x^{2}\]

OpenStudy (zarkon):

no

OpenStudy (zarkon):

du=-dx

OpenStudy (zarkon):

u=1-x so x=1-u and \(x^2=(1-u)^2=1-2u-u^2\)

OpenStudy (anonymous):

so then (1-2u-u^2)du?

OpenStudy (zarkon):

no

OpenStudy (zarkon):

\[x^2=1-2u-u^2\] \[1-x=u\] \[dx=-du\] combine

OpenStudy (zarkon):

1-x=u \[\sqrt{1-x}=\sqrt{u}\]

OpenStudy (anonymous):

ok, i see there is (1-2u-u^2)(u^1/2)(du), if am correct,

OpenStudy (anonymous):

typo there: (1-2u+u^2) not -u^2

OpenStudy (zarkon):

yes..now distribute the \(u^{1/2}\) yes...typo from me above too

OpenStudy (zarkon):

you are still missing the negative sign from the dx=-du

OpenStudy (anonymous):

ah, yes missed that

OpenStudy (zarkon):

\[\int x^2(1-x)^{1/2}dx\] \[=\int (1-2u+u^2)(u^{1/2})(-du)\]

OpenStudy (anonymous):

thank you for the help, i figured the question, and by the way data is my favorite from star trek next gen

OpenStudy (zarkon):

good, because he is awesome ;)

OpenStudy (anonymous):

lol